# Qubit-assisted quantum metrology

### v1

@@ IV Quantum Fisher information
Suppose that we have found an exact solution about $T$ on demand of the time-reversal strategy, then the whole unitary evolution operator in Eq. (2) can be rewritten as
@@
Using the Baker-Campbell-Hausdorff formula, we have (the details are provided in Appendix A)
@@
where the phase generator forms as $J(\phi)\equiv\cos(\phi)J_{x}+\sin(\phi)J_{y}$ with $\phi=\omega_{P}t_{1}$. We are now on the stage of analyzing the precision limit of the qubit-assisted metrology.
Initially, the ancillary qubit is assumed to be a pure qubit $\rho_{A}=|\varphi\rangle\langle\varphi|$ with Kitagawa and Ueda (1993); Jin et al. (2009)
@@
Here $\theta_{0}$ and $\phi_{0}$ determine the population imbalance and the relative phase between the two bases, respectively. Without loss of generality, we set the azimuthal angle $\phi_{0}=0$. For the input state of the probe system, always it can be written as
@@
Using Eqs. (14) and (15), QFI of the output state associated with the metrology precision for estimating $\theta$ is given by Braunstein and Caves (1994); Braunstein et al. (1996); Zhang et al. (2013); Liu et al. (2014)
@@
with $p_{i}\neq 0$. We actually obtain a difference between a positive expression, that is vital to achieve Heisenberg scaling as shown in the following, and a non-positive one. A straightforward method to enhance QFI is thus to minimize the magnitude of the second term in Eq. (16) or more generally and precisely to minimize the expression inside the absolute value. Using Eq. (13), we have
$\langle\psi_{i}|\langle\varphi|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\varphi\rangle|\psi_{i}\rangle=\cos(gt_{1})\langle J(-\phi)\rangle_{\psi_{i}}$ (17)
$-\sin(gt_{1})\langle\sigma_{z}\rangle_{\varphi}\left\langle J\left(\frac{\pi}{2}-\phi\right)\right\rangle_{\psi_{i}},$
where $\langle A\rangle_{\psi}\equiv\langle\psi|A|\psi\rangle$. Under a proper state of the ancillary qubit and an optimized evolution time for Step $1$, i.e.,
@@
with an integer $n$, the two terms in Eq. (17) can vanish at the same time for arbitrary probe state $\psi_{i}$. Consequently, Eq. (13) becomes $U_{\theta}=\exp[i\theta J(\pi/2-\phi_{\rm opt})\sigma_{z}]$ with $\phi_{\rm opt}\equiv\omega_{P}t_{1,{\rm opt}}$.
We emphasize that Eq. (17) presents the main advantage of our qubit-assisted metrology over a conventional parameter estimation Giovannetti et al. (2011); Tóth and Apellaniz (2014); Nawrocki (2019); Polino et al. (2020). For a conventional metrology, in which only the probe state is considered, Eq. (17) becomes
$\langle\psi_{i}|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\psi_{i}\rangle=\langle\cos(\omega_{P}t_{1})J_{x}-\sin(\omega_{P}t_{1})J_{y}\rangle_{\psi_{i}}.$ (19)
Hardly one can find a state-independent time $t_{1}$ leading to $\langle\psi_{i}|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\psi_{i}\rangle=0$, since $\cos(\omega_{P}t_{1})$ and $\sin(\omega_{P}t_{1})$ cannot simultaneously become zero.
Figure 3: (a) QFI as a function of $\theta_{0}$ with a fixed $t_{1}=\pi/(2g)$ and various probe-spin number $N$, (b) QFI as a function of the renormalized evolution time $gt_{1}$ with a fixed $\theta_{0}=\pi/2$ and various $N$, (c) Renormalized QFI $F_{Q}/N^{2}$ in space of $\theta_{0}$ and $gt_{1}$ with $N=4$, and (d) QFI as a function of $N$ for $\rho_{P}=\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$, $\theta_{0}=\pi/2$, and $t_{1}=\pi/(2g)$ (red crosses); for $\rho_{P}^{\rm opt}$, $\theta_{0}=\pi/2$, and $t_{1}=3\pi/(4g)$ (orange squares); for $\rho_{P}^{\rm opt}$, $\theta_{0}=\pi/8$, and $t_{1}=\pi/(2g)$ (green triangles); for $\rho_{P}^{\rm th}$ with $\beta=1$, $\theta_{0}=\pi/2$, and $t_{1}=\pi/(2g)$ [blue circles for numerical results and blue solid line for analytical expression in Eq. (23)]. In (a), (b), (c), and (d), the black dashed line and the black dot-dashed line indicate the Heisenberg scaling $N^{2}$ and the shot-noise scaling $N$, respectively. $\omega_{P}=\omega_{A}=3g$
@@
$F_{Q}=4\sum_{i=1}^{d}p_{i}\left\langle J\left(\frac{\pi}{2}-\phi_{\rm opt}\right)^{2}\right\rangle_{\psi_{i}},$ (20)
@@
which is a weighted average over the expectation value for the operator $J(\pi/2-\phi_{\rm opt})^{2}$ within each eigenstate $\{|\psi_{i}\rangle\}$ of $\rho_{P}$. Straightforwardly, QFI can attains its peak value $F_{Q}=N^{2}$ when the probe spin ensemble is prepared in a pure state $|j,j\rangle_{\rm opt}$, or $|j,-j\rangle_{\rm opt}$, and or a mixed state over them, such as $\rho_{P}=(|j,j\rangle_{\rm opt}\langle j,j|+|j,-j\rangle_{\rm opt}\langle j,-j|)/2$, where $|j,m\rangle_{\rm opt}$ with $-j\leq m\leq j$ denotes the eigenstate for the optimized collective angular momentum operator $J(\pi/2-\phi_{\rm opt})$. Subjected to the duration time $t_{1,{\rm opt}}$ and the time-reversal strategy, the bases $|j,m\rangle_{\rm opt}$’s are determined by the magnitudes of $g$, $\omega_{P}$, and $\omega_{A}$. Heisenberg scaling can therefore be approached under optimized $\theta_{0}$ and $t_{1}$.
@@
This condition could be confirmed by the numerical simulation in Fig. 3, where the input state of the probe ensemble is set as $\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$ and the frequencies of both probe spin and ancillary qubit are the same as in Fig. 2. Figure 3(a) describes the dependence of the quantum Fisher information $F_{Q}$ on the input parameter of the ancillary qubit $\theta_{0}$ under an optimized evolution time $t_{1,{\rm opt}}$ given by Eq. (18). It is shown that $F_{Q}$ follows the Heisenberg scaling when $\theta_{0}=\pi/2$ for various $N$. In Fig. 3(b), the initial state of the ancillary qubit is fixed with $\theta_{0}=\pi/2$ and $F_{Q}$ finds its peak at $t_{1}=\pi/(2g)$ as expected by Eq. (18). In both Fig. 3(a) and Fig. 3(b), $F_{Q}$ is symmetrical to the optimized points and varies smoothly around them, indicating that our metrology protocol is not sensitive to the imperfections in parametric control.
@@
In Fig. 3(c), we take $N=4$ and demonstrate the renormalized QFI $F_{Q}/N^{2}$ in the space of $\theta_{0}$ and $t_{1}$. The black dot-dashed lines are the contour lines indicating the proportional coefficients. The central region around point-$A$ describes the most optimized condition that is in agreement with Eq. (18). The surrounding regions about point-$B$ and point-$C$ describe less optimized conditions with $t_{1}$ or $\theta_{0}$ largely departing from Eq. (18). The behaviors for these points with various $N$ are plotted with the red crosses, the orange squares, and the green triangles, respectively, in Fig. 3(d). The first one sticks to its upper bound $F_{Q}=N^{2}$, and the last two are apparently below this bound. However, a similar behavior to Heisenberg scaling $F_{Q}\propto N^{2}$ can be restored when $N$ is sufficiently large, although $\theta_{0}$ or $gt_{1}$ is not optimized.
@@
It is interesting to find that QFI could approach Heisenberg scaling even if the probe ensemble is prepared as a thermal state in the bases of $|j,m\rangle_{\rm opt}$, i.e.,
$\rho_{P}^{\rm th}=\frac{1}{Z_{\beta}}e^{-\beta J(\pi/2-\phi_{\rm opt})},$ (21)
where $Z_{\beta}={\rm Tr}\{\exp[-\beta J(\pi/2-\phi_{\rm opt})]\}$ is the partition function and $\beta$ is the inverse temperature. By Eqs. (20) and (21), we have
@@
And then for a large probe-spin number, i.e., $N\gg 1$, it is approximated as
@@
Clearly $F_{Q}=N^{2}$ when $\beta\rightarrow\infty$. It is indicated that the Heisenberg-scaling behavior $N^{2}$ dominates QFI at least in the low-temperature limit. And $F_{Q}$ will become divergent in the high-temperature limit, i.e., $\beta\rightarrow 0$, which cannot be used to estimate the metrology precision. In Fig. 3(d), the analytical expression in Eq. (23) is confirmed to match the numerical result and both of them approach the Heisenberg scaling, provided $N\gg 1$.
@@ V Classical Fisher information
In a practical scenario of parametric estimation, CFI is defined by catching the amount of information encoded in the probability distribution for the output state under the estimation process Braun et al. (2018); Liu et al. (2020); Tan et al. (2021). And it is upper bounded by its quantum counterpart on the metrology precision. With the time-reversal strategy, it was shown Yurke et al. (1986); Agarwal and Davidovich (2022); Wang et al. (2024) that CFI can be saturated with QFI, which indicates that all the information encoded in the probe state from the output probability distribution has been extracted. We now derive CFI of our protocol under the same optimized settings in Eq. (18) with $n=0$ as for QFI. For the input state $\rho_{P}\otimes\rho_{A}$ indicated by Eqs. (14) and (15), the output state after the two-step evoluation reads
@@
where $p_{i}$’s are the eigenvalues for the eigenstates $|\psi_{i}\rangle$ of $\rho_{P}$ and $p_{i,m,m^{\prime}}\equiv\langle j,m|\psi_{i}\rangle\langle\psi_{i}|j,m^{\prime}\rangle_{\rm opt}$. Subsequently, we perform the projective measurements $\{|\Psi_{m,\pm}\rangle\langle\Psi_{m,\pm}|\}$ on the full system as described in Fig. 1, where $|\Psi_{m,\pm}\rangle=|j,m\rangle_{\rm opt}\otimes|\pm\rangle$ with |±$|\pm\rangle=(|e\rangle\pm|g\rangle)/\sqrt{2}$, the eigenstates of $\sigma_{x}$ for the ancillary qubit. Note $|j,m\rangle_{\rm opt}$’s are eigenstates of $J(\pi/2-\phi_{\rm opt})$ determined by $g$, $\omega_{P}$, and $\omega_{A}$ in advance. The probability about detecting the output state in the basis $|\Psi_{m,\pm}\rangle$ reads
@@
$=\frac{1\pm\cos(2m\theta)}{2}\sum_{i=1}^{d}p_{i}p_{i,m,m}.$
By the probability distribution $P(m,\pm|\theta)$ Braun et al. (2018); Liu et al. (2020); Tan et al. (2021), CFI can be calculated as
@@
$=$ $4\sum_{i=1}^{d}p_{i}\langle\psi_{i}|J\left(\frac{\pi}{2}-\phi_{\rm opt}\right)^{2}|\psi_{i}\rangle=F_{Q}.$
The last line is exactly the same as Eq. (20) under the optimized condition in Eq. (18), independent of $\rho_{P}$. It is thus verified that QFI can attain its upper-bound in our qubit-assisted metrology. In addition, when the probe is prepared as a polarized state $\rho_{P}^{\rm opt}=|j,\pm j\rangle_{\rm opt}\langle j,\pm j|$, Eq. (25) reduces to
@@
which implies that the magnitude of $\theta$ can be inferred from the probability $P(j,+|\theta)$ and $P(j,-|\theta)$. In this case, only performing the projective measurement on the ancillary qubit $\rho_{A}$ yields the Heisenberg-scaling limit $F_{c}=F_{Q}=N^{2}$.
Figure 4: (a) Renormalized CFI $F_{c}/N^{2}$ in space of $gt_{1}$ and $gt_{2}$. (b) Renormalized CFI $F_{c}/N^{2}$ as a function of $gt_{2}$ with a fixed $gt_{1}=\pi/2$ (blue solid line) and its maximum value obtained with a varying $gt_{1}$ (red dashed line). The vertical black dashed line represents the analytical result in Eq. (9). The probe spin ensemble and the ancillary qubit are initialized in the state $\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$ and $|\varphi\rangle=(|e\rangle+|g\rangle)/\sqrt{2}$, respectively. The other parameters are $\phi_{\rm opt}=\omega_{P}\pi/(2g)$, $N=5$, and $\omega_{P}=\omega_{A}=3g$.
@@
It is interesting to find that CFI also attains the Heisenberg scaling $F_{c}=N^{2}$ when $t_{2}=\pi/(2g)$ [see the vertical black dot-dashed line in Fig. 4(b)]. It is beyond the sufficient condition for the time-reversal strategy as illustrated by Eq. (9) and Fig. 2. However, it is straightforward to see that when $t_{1}=t_{2}=\pi/(2g)$, the full time-evolution operator in the end can be written as $U^{\prime}_{\theta}=U(\pi/g)U^{\dagger}(t_{1}){R}_{x}(\theta)U(t_{1})=\sigma_{z}U_{\theta}$, where $U_{\theta}$ is given by Eq. (13). Then using Eq. (24), we have $\rho^{\prime}(\theta)=\sigma_{z}U_{\theta}\rho_{P}\otimes\rho_{A}U_{\theta}^{\dagger}\sigma_{z}=\sigma_{z}\rho(\theta)\sigma_{z}$ and hence the probability distribution in Eq. (29) becomes $P^{\prime}(j,\pm|\theta)=[1\mp\cos(2m\theta)]/2$, which means $P^{\prime}(j,\pm|\theta)=P(j,\mp|\theta)$. It indicates that one can still extract all the information encoded in the probe state from the output probabilities and thus $F_{c}=N^{2}$.
VI Discussion
Quantum metrology is conventionally regarded as a parametric-estimation protocol using entanglement or quadratical squeezing to improve the estimation precision beyond the standard quantum limit with respect to the system size $N$. A benchmark result is that the photonic entangled NOON states are capable to attain the Heisenberg scaling characterized with an $N^{2}$-dependent quantum Fisher information. However, a feasible implementation is limited to the number of photons or excitations due to the difficulty in generating a large-scale and high-fidelity NOON state Bollinger et al. (1996); Walther et al. (2004); Afek et al. (2010); Zhang et al. (2013), which achieves up to $N=5$ so far in experiment Afek et al. (2010). As for the spin squeezing that is generated in the nonlinear atom-atom interaction platform, the OAT squeezing demonstrates a scaling below the Heisenberg-limited measurement precision with $\propto 1/N^{2/3}$, e.g., in the cold thermal ensembles Hosten et al. (2016), the metrological gain is then found to be up to $(\delta\theta_{\rm SQL}/\delta\theta)^{2}\sim 100$ with $10^{6}$ atoms.
In contrast, our qubit-assisted metrology protocol can be enhance QFI up to the Heisenberg scaling of $N^{2}$ with respect to the total probe-spin number $N$, even when the probe system is in a mixed state. More than the $ZZ$ interaction between the ancillary qubit and the probe large-spin in the full Hamiltonian (1), our protocol is applicable to another spin-spin-bath model Caldeira et al. (1993); Shao et al. (1996); Xu et al. (2011); Wang and Shao (2012) with Hamiltonian
@@ VI Discussion
The corresponding circuit model for the qubit-assisted metrology in Fig. 1 is almost invariant, except that the encoded phase parameter $\theta$ is now obtained by another rotation $R_{z}(\theta)=\exp(-i\theta J_{z})$ of the probe system. Thus the circuit is described by
@@
Using the modified full Hamiltonian (30), Eq. (5) becomes
@@
where $\tilde{C}_{e/g,m}=\exp[-i(m\tilde{\omega}\pm\omega_{A})T]$ with $\tilde{\omega}\equiv\sqrt{\omega^{2}_{P}+g^{2}}$. Here $|j,m\rangle_{\phi}$ and $|j,m^{\prime}\rangle_{\phi^{\prime}}$ with eigenvalues $m$ and $m^{\prime}$ denote the eigenstates of the operators $(\omega_{P}J_{z}+gJ_{x})/\tilde{\omega}$ and $(\omega_{P}J_{z}-gJ_{x})/\tilde{\omega}$, respectively. According to $U_{\theta}$ in Eq. (31) and the time-reversal condition in Eq. (3), we have
@@
A sufficient solution for Eq. (34) is found to be independent of the parity of the probe-spin number $N$, and reads,
@@
where $n_{j}$’s, $j=5,6$, are proper integers.
With a proper $T$ on the time-reversal strategy, the whole unitary evolution operator in Eq. (31) can be rewritten as
@@
Using the Baker-Campbell-Hausdorff formula, we have

### v2

@@ IV Quantum Fisher information
Suppose that we found an exact solution about $T$ on demand of the time-reversal strategy, then the entire unitary evolution operator in Eq. (2) can be rewritten as
@@
Using the Baker-Campbell-Hausdorff (BCH) formula, we have (the details are provided in Appendix A)
@@
where the phase generator is defined as $J(\phi)\equiv\cos(\phi)J_{x}+\sin(\phi)J_{y}$ with $\phi=\omega_{P}t_{1}$. We are now on the stage of analyzing the precision limit of the qubit-assisted metrology.
Initially, the ancillary qubit is assumed to be a pure state $\rho_{A}=|\varphi\rangle\langle\varphi|$ with Kitagawa and Ueda (1993); Jin et al. (2009)
@@
Here $\theta_{0}$ and $\phi_{0}$ determine the population imbalance and the relative phase between the two bases, respectively. Without loss of generality, we set the azimuthal angle $\phi_{0}=0$. For the input state of the probe system, it always can be written as
@@
Using Eqs. (14) and (15), QFI of the output state for estimating $\theta$ is given by Braunstein and Caves (1994); Braunstein et al. (1996); Zhang et al. (2013); Liu et al. (2014)
@@
with $p_{i}\neq 0$. It is a difference between two positive terms. The first term is crucial to achieve Heisenberg scaling as shown in the following. A straightforward idea to enhance QFI is thus to minimize the second term, or more generally and precisely, to minimize the expression magnitude inside the absolute value. Using Eq. (13), we have
$\langle\psi_{i}|\langle\varphi|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\varphi\rangle|\psi_{j}\rangle=-i\cos(gt_{1})\langle\psi_{i}|J(-\phi)|\psi_{j}\rangle$ (17)
$+i\sin(gt_{1})\langle\sigma_{z}\rangle_{\varphi}\langle\psi_{i}|J\left(\frac{\pi}{2}-\phi\right)|\psi_{j}\rangle,$
where $\langle\sigma_{z}\rangle_{\varphi}=\langle\varphi|\sigma_{z}|\varphi\rangle$. Under a proper state of the ancillary qubit and an optimized joint-evolution time for Step $1$, i.e.,
@@
with $n$ integer, the two terms in Eq. (17) can vanish at the same time for arbitrary probe state $|\psi_{i}\rangle$. Consequently, Eq. (13) becomes $U_{\theta}=\exp(i\theta J_{\rm opt}\sigma_{z})$ with $J_{\rm opt}\equiv J(\pi/2-\phi_{\rm opt})$ and $\phi_{\rm opt}\equiv\omega_{P}t_{1,{\rm opt}}$.
Equation (17) indicates the main advantage of our qubit-assisted metrology over a conventional parameter estimation Giovannetti et al. (2011); Tóth and Apellaniz (2014); Nawrocki (2019); Polino et al. (2020). The presence of the ancillary qubit provides an extra degree of freedom for control. In the absence of the ancillary qubit, i.e., the coupling strength between the probe system and the ancillary qubit $g=0$, Eq. (16) becomes
$F_{Q}$ $=\sum_{i=1}^{d}4p_{i}\langle\psi_{i}|J^{2}(-\phi)|\psi_{i}\rangle$ (19)
$-\sum_{i,j=1}^{d}\frac{8p_{i}p_{j}}{p_{i}+p_{j}}|\langle\psi_{i}|J(-\phi)|\psi_{j}\rangle|^{2},$
where the second term is determined by
$\langle\psi_{i}|J(-\phi)|\psi_{j}\rangle=\langle\psi_{i}|\cos(\omega_{P}t_{1})J_{x}-\sin(\omega_{P}t_{1})J_{y}|\psi_{j}\rangle.$ (20)
In contrast to Eq. (17), one can hardly find a state-independent time $t_{1}$ leading to $\langle\psi_{i}|J(-\phi)|\psi_{j}\rangle=0$, simply because $\cos(\omega_{P}t_{1})$ and $\sin(\omega_{P}t_{1})$ cannot simultaneously become zero. The only exception is when the probe is initialized as a GHZ state, i.e., $|\psi\rangle=(|j,j\rangle_{\phi}+|j,-j\rangle_{\phi})/\sqrt{2}$. Here $|j,m\rangle_{\phi}$ with $-j\leq m\leq j$ are the eigenstates of the collective angular momentum operator $J(-\phi)$.
Figure 3: (a) QFI as a function of $\theta_{0}$ with a fixed $gt_{1}=\pi/2$ and various probe-spin number $N$, (b) QFI as a function of the renormalized evolution time $gt_{1}$ with a fixed $\theta_{0}=\pi/2$ and various $N$, (c) Renormalized QFI $F_{Q}/N^{2}$ in space of $\theta_{0}$ and $gt_{1}$ with $N=4$, and (d) QFI as a function of $N$ for different probe states and parameters. The red crosses indicate point A in (c) for $\rho_{P}=\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$ with $\theta_{0}=\pi/2$ and $gt_{1}=\pi/2$. The orange squares indicate point B for $\rho_{P}^{\rm opt}$ with $\theta_{0}=\pi/2$ and $gt_{1}=3\pi/4$. The green triangles denote point C for $\rho_{P}^{\rm opt}$ with $\theta_{0}=\pi/8$ and $gt_{1}=\pi/2$. The blue circles (labeled by D) and the blue solid line are numerical simulation and analytical results for thermal state $\rho_{P}=\rho_{P}^{\rm th}$ with $\beta=1$, $\theta_{0}=\pi/2$, and $gt_{1}=\pi/2$, respectively. In (a), (b), and (d), the black dashed line indicates the Heisenberg scaling and in (d), the black dot-dashed line indicates the shot-noise scaling. $\omega_{P}=\omega_{A}=3g$.
@@
$F_{Q}=4\sum_{i=1}^{d}p_{i}\langle J_{\rm opt}^{2}\rangle_{\psi_{i}},$ (21)
@@
which is a weighted average over the expectation value for the operator $J_{\rm opt}^{2}$ in each eigenstate $|\psi_{i}\rangle$ of $\rho_{P}$. The peak value for QFI $F_{Q}=N^{2}$ is attainable when the probe system is prepared in a pure state $|j,j\rangle_{\rm opt}$, $|j,-j\rangle_{\rm opt}$, or an arbitrary mixed or superposed state over them, such as $(|j,j\rangle_{\rm opt}\langle j,j|+|j,-j\rangle_{\rm opt}\langle j,-j|)/2$ or $(|j,j\rangle_{\rm opt}+|j,-j\rangle_{\rm opt})/\sqrt{2}$. Here $|j,m\rangle_{\rm opt}$ with $-j\leq m\leq j$ denote the eigenstates of the optimized collective angular momentum operator $J_{\rm opt}$. Subject to the duration time $t_{1,{\rm opt}}$ and the time-reversal strategy, the bases $|j,m\rangle_{\rm opt}$’s are determined by the magnitudes of $g$, $\omega_{P}$, and $\omega_{A}$. Heisenberg scaling can therefore be attainable under optimized $\theta_{0}$ and $t_{1}$.
@@
This condition could be confirmed by the numerical simulation in Fig. 3, where the input state of the spin ensemble is set as $\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$ if not otherwise stated. Figure 3(a) describes the dependence of the quantum Fisher information $F_{Q}$ on the input parameter of the ancillary qubit $\theta_{0}$ under an optimized evolution time $t_{1,{\rm opt}}$ in Eq. (18). It is shown that $F_{Q}$ follows the Heisenberg scaling when $\theta_{0}=\pi/2$ for various $N$. In Fig. 3(b), the initial state of the ancillary qubit is fixed with $\theta_{0}=\pi/2$ and $F_{Q}$ arrives at its peak value $F_{Q}=N^{2}$ at $gt_{1}=\pi/2$ as expected by Eq. (18).
In both Figs. 3(a) and  3(b), $F_{Q}$ is symmetrical to the optimized points and varies smoothly around them, indicating that our metrology protocol is not sensitive to the deviation in parametric control. If $g\rightarrow g^{\prime}=g+\Delta g$ and $\omega_{P}\rightarrow\omega^{\prime}_{P}=\omega_{P}+\Delta\omega_{P}$ in Eq. (16), where $|\Delta gt_{1,{\rm opt}}|\ll 1$ and $|\Delta\omega_{P}t_{1,{\rm opt}}|\ll 1$ indicate the relative deviations in system parameters, then QFI using the optimized probe state, e.g., $\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$, becomes
$F_{Q}\approx N^{2}-N(N-1)(\Delta\omega_{P}^{2}+\Delta g^{2})t_{1,{\rm opt}}^{2},$ (22)
up to the second order of $\Delta g$ and $\Delta\omega_{P}$. The details are provided in Appendix B. This result again confirms that our condition in Eq. (18) is optimal.
@@
In Fig. 3(c), we take $N=4$ and demonstrate the renormalized QFI $F_{Q}/N^{2}$ in the space of $\theta_{0}$ and $t_{1}$. The black dot-dashed lines are the contour lines. The central region around point $A$ describes the most optimized condition that is in agreement with Eq. (18). The surrounding regions about point $B$ and point $C$ describe the sub-optimized conditions with $t_{1}$ or $\theta_{0}$ largely departing from Eq. (18). The behaviors for these points with various $N$ are plotted with the red crosses, the orange squares, and the green triangles, respectively, in Fig. 3(d). The first one sticks to the upper bound $F_{Q}=N^{2}$. Although the rest two are apparently below this bound, an asymptotic behavior to Heisenberg scaling $F_{Q}\propto N^{2}$ can appear when $N$ is sufficiently large.
@@
It is interesting to find that QFI could approach the Heisenberg scaling even if the probe ensemble is degraded to be a thermal state in the bases of $|j,m\rangle_{\rm opt}$, i.e.,
$\rho_{P}^{\rm th}=\frac{1}{Z_{\beta}}e^{-\beta J_{\rm opt}},$ (23)
where $Z_{\beta}={\rm Tr}[\exp(-\beta J_{\rm opt})]$ is the partition function and $\beta$ is the inverse temperature. By Eqs. (21) and (23), we have
@@
Then for a large-size probe, i.e., $N\gg 1$, it is approximated as
@@
Clearly, $F_{Q}=N^{2}$ when $\beta\rightarrow\infty$. It indicates that the Heisenberg-scaling behavior dominates QFI at least in the low-temperature limit. Despite $F_{Q}$ will become divergent in the high-temperature limit, i.e., $\beta\rightarrow 0$, it cannot be used to estimate the metrology precision. The analytical expression in Eq. (25) is confirmed to match the numerical result in Fig. 3(d) and both of them (see the blue solid line and the blue circles labeled by $D$) approach the Heisenberg scaling for a sufficient large $N$.
In sharp contrast to the conventional protocol with no ancillary qubit [see, e.g., Eq. (19)], our metrology protocol is much broader in optimized probe states. Beyond the GHZ state in an appropriate basis, e.g., $|j,m\rangle_{\phi}$, the Heisenberg limit in metrology precision can be achieved by a product state or even a finite-temperature state.
@@ V Classical Fisher information
In a practical scenario of parametric estimation, CFI is defined by the amount of information encoded in the probability distribution for the output state Braun et al. (2018); Liu et al. (2020); Tan et al. (2021). In addition, it is upper bounded by its quantum counterpart on the metrology precision. With the time-reversal strategy, it was shown Yurke et al. (1986); Agarwal and Davidovich (2022); Wang et al. (2024) that CFI can be saturated with QFI, indicating that all the information encoded in the output probability distribution has been extracted. We here derive the CFI of our protocol under the same optimized settings in Eq. (18) with $n=0$ as for QFI. For the input state $\rho_{P}\otimes\rho_{A}$ indicated by Eqs. (14) and (15), the output state after the two-step evolution (13) reads
@@
where $p_{i}$’s are the weights of the eigenstates $|\psi_{i}\rangle$ in $\rho_{P}$ and $p_{i,m,m^{\prime}}\equiv\langle j,m|\psi_{i}\rangle\langle\psi_{i}|j,m^{\prime}\rangle_{\rm opt}$. Subsequently, we perform the projective measurements $|\Psi_{m,\pm}\rangle\langle\Psi_{m,\pm}|$ on the full system as described in Fig. 1, where $|\Psi_{m,\pm}\rangle=|j,m\rangle_{\rm opt}\otimes|\pm\rangle$ with |±$|\pm\rangle=(|e\rangle\pm|g\rangle)/\sqrt{2}$, the eigenstates of $\sigma_{x}$ for the ancillary qubit. The probability of detecting the output state in $|\Psi_{m,\pm}\rangle$ reads
@@
$=\frac{1\pm\cos(2m\theta)}{2}\sum_{i=1}^{d}p_{i}p_{i,m,m},$
by which CFI Braun et al. (2018); Liu et al. (2020); Tan et al. (2021) can be calculated as
@@
$=$ $4\sum_{i=1}^{d}p_{i}\langle\psi_{i}|J_{\rm opt}^{2}|\psi_{i}\rangle=F_{Q}.$
The last line is exactly the same as Eq. (21) under the optimized condition in Eq. (18), irrespective of the probe state $\rho_{P}$. It is thus verified that CFI can attain its upper-bound in our metrology. In addition, when the probe is polarized as $\rho_{P}^{\rm opt}=|j,\pm j\rangle_{\rm opt}\langle j,\pm j|$, Eq. (27) reduces to
@@
which implies that the magnitude of $\theta$ can be inferred from the probabilities $P(j,+|\theta)$ and $P(j,-|\theta)$. In this case, only performing the projective measurement on the ancillary qubit can yield the Heisenberg-scaling limit $F_{c}=F_{Q}=N^{2}$.
Figure 4: (a) Renormalized CFI $F_{c}/N^{2}$ in space of $gt_{1}$ and $gt_{2}$. (b) Dynamics of the renormalized CFI $F_{c}/N^{2}$ versus $gt_{2}$ with a fixed $gt_{1}=\pi/2$ (blue solid line) and its maximum value obtained with a varying $gt_{1}$ (red dashed line). The vertical black dashed line represents the analytical result in Eq. (9). The probe-spin ensemble and the ancillary qubit are initialized in the state $\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$ and $|\varphi\rangle=(|e\rangle+|g\rangle)/\sqrt{2}$, respectively. The other parameters are $\phi_{\rm opt}=\omega_{P}\pi/(2g)$, $N=5$, and $\omega_{P}=\omega_{A}=3g$.
@@
CFI can also attain the Heisenberg scaling $F_{c}=N^{2}$ when $gt_{2}=\pi/2$ [see the vertical black dot-dashed line in Fig. 4(b)]. It is not covered by the sufficient condition for the time-reversal strategy as illustrated by Eq. (9) and Fig. 2. However, it is straightforward to see that with $gt_{1}=gt_{2}=\pi/2$, the entire evolution operator can be written as $U^{\prime}_{\theta}=U(\pi/g)U^{\dagger}(t_{1}){R}_{x}(\theta)U(t_{1})=\sigma_{z}U_{\theta}$, where $U_{\theta}$ is given by Eq. (13). Then by using Eq. (26), we have $\rho^{\prime}(\theta)=\sigma_{z}U_{\theta}\rho_{P}\otimes\rho_{A}U_{\theta}^{\dagger}\sigma_{z}=\sigma_{z}\rho(\theta)\sigma_{z}$ and hence the probability distribution in Eq. (31) becomes $P^{\prime}(j,\pm|\theta)=[1\mp\cos(2m\theta)]/2$, which means $P^{\prime}(j,\pm|\theta)=P(j,\mp|\theta)$. It indicates that still one can extract all the information encoded in the probe state from the output probabilities and thus $F_{c}=N^{2}$.
VI metrology with $XZ$ interaction
The preceding theoretical framework does not confine to the $ZZ$ interaction in the Hamiltonian (1). Our protocol is found to be applicable to another typical spin-spin-bath model Caldeira et al. (1993); Shao et al. (1996); Xu et al. (2011); Wang and Shao (2012) with a Hamiltonian
@@
The relevant circuit model for our metrology in Fig. 1 is almost invariant, except that the encoded phase parameter $\theta$ is now obtained by another rotation $R_{z}(\theta)=\exp(-i\theta J_{z})$ of the probe system. Thus the entire unitary evolution is described by
@@
By using the modified full Hamiltonian (32), Eq. (5) becomes
@@
where $\tilde{C}_{e/g,m}=\exp[-i(m\tilde{\omega}\pm\omega_{A})T]$ with $\tilde{\omega}\equiv\sqrt{\omega^{2}_{P}+g^{2}}$. Here $|j,m\rangle_{\phi}$ and $|j,m^{\prime}\rangle_{\phi^{\prime}}$ denote the eigenstates of the operators $(\omega_{P}J_{z}+gJ_{x})/\tilde{\omega}$ and $(\omega_{P}J_{z}-gJ_{x})/\tilde{\omega}$ and the relevant eigenvalues are $m$ and $m^{\prime}$, respectively. According to $U_{\theta}$ in Eq. (33) and the time-reversal condition in Eq. (3), we have
@@
A sufficient solution for Eq. (36) reads,
@@
where $n_{j}$’s, $j=5,6$, are proper integers. It is independent of the parity of the probe-spin number $N$.
With a proper $T$ on the time-reversal strategy, Eq. (33) can be rewritten as
@@
By using the BCH formula, we have